OpenStudy (anonymous):

If cos θ=–5/13 and sin θ < 0, what is tan θ

OpenStudy (anonymous):

\[\sin \theta = - \sqrt{1 - \cos ^{2} \theta} \]

OpenStudy (anonymous):

solve for sin theta

OpenStudy (anonymous):

i need to know what is tan 0

OpenStudy (anonymous):

If sine is negative and tangent is positive, then cosine is ____?____ tangent=sine/cosine=negative/cosine=positive =>cosine is ___?____ Is it positive or negitive?

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

Right. You want tangent to be positive not negative. So you need cosine to be negative in order for tangent to be positive since sine is negative.

OpenStudy (anonymous):

alright then so what is tan 0

OpenStudy (anonymous):

sin ( 0 ) = -5/13 is given to you

zepdrix (zepdrix):

This triangle satisfies \(\Large\rm \cos\theta=-\frac{5}{13}\). You can use your Pythagorean Theorem to find the missing side :)

zepdrix (zepdrix):

Sine was not given to us Sarah D: Maybe you misread it.

OpenStudy (anonymous):

i dont understand what is tan 0

zepdrix (zepdrix):

Tangent is `opposite` over `adjacent`, yes? So we first need to find the length of the `opposite` side.`

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

So use your Pythagorean Theorem: \(\Large\rm (-5)^2+(opposite)^2=13^2\)